on an eight question true -false quiz a student guesses at each answer without even reading the question. what is the probability that he gets at least one of the answers correct
Answer by swincher4391(1107)

Whenever you see the word "at least" think about the complement. It is much easier to calculate the probability of "none" than "all other possibilities". What is the probability that you get no answers right at all?
P[none]= (1/2)^8
So P[at least correct] = 1 - P[none correct] = 1 - (1/2)^8 = .9961